OpenStudy (loser66):

find a, b satisfy \(\lim_{x\rightarrow 0}\dfrac{\sqrt{ax+b}-2}{x}=1\) Please, help

OpenStudy (loser66):

@agent0smith

OpenStudy (agent0smith):

The numerator and denominator must have the same value, for it to equal 1. So set them equal to each other, and plug in x=0.

OpenStudy (loser66):

How can we say so? it is a limit, not a fraction.

OpenStudy (loser66):

\(f'(0) =\lim_{x\rightarrow 0}\dfrac{f(x+0) -f(0)}{x-0}=1\) So, \(f(x) =\sqrt{ax+b}\), \(f(0) =2, f'(0) =1\) \(f'(x) =\dfrac{a}{2\sqrt{ax+b}}\) \(f'(0) = \dfrac{a}{2b}=1\), so a =2b \(f(0) =2 =\sqrt b\), Hence, b=4, a=8

OpenStudy (loser66):

However, when I plug back to get the lim, I got it wrong.

imqwerty (imqwerty):

*\(f'(0) = \dfrac{a}{2\color{red}{\sqrt{b}}}=1\)

OpenStudy (loser66):

Bingo!! thank you

OpenStudy (loser66):

Now, everything go right!!

imqwerty (imqwerty):

=]

OpenStudy (agent0smith):

What I said works for getting b. If the value of the numerator is not 0, then the limit would be infinity.

OpenStudy (agent0smith):

Since \[\large \dfrac{\sqrt{ax+b}-2}{x}\] must be either of the form \(\infty/\infty\) or 0/0 for it to be equal to 1. And we know it's not \(\infty/\infty\)...

OpenStudy (loser66):

Got you. Thank you so much.